OpenStudy (anonymous):

how do you turn this into standard form: -(x-2)^2 +1

OpenStudy (anonymous):

multiply out and combine like terms start with \[-(x-2)(x-2)+1\] and go from there

OpenStudy (anonymous):

or you might remember that \(9a+b)^2=a^2+2ab+b^2\) so \[(x-2)^2=x^2-4x+4\] giving \[-(x^2-4x+4)+1\] then distribute the minus sign, then add 1

OpenStudy (anonymous):

Thanks for answering my question! So, would the answer be -x^2 + 4x - 3 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks!!! And to take a function in vertex form and turn it into standard form, would you use the same method?

OpenStudy (anonymous):

it was in vertex form

OpenStudy (anonymous):

vertex form is \[y=a(x-h)^2+k\]

OpenStudy (anonymous):

how would I turn the standard form back into vertex form ?

OpenStudy (anonymous):

a) complete the square or b) find the vertex

OpenStudy (anonymous):

Ohh okay thanks!!

OpenStudy (anonymous):

yw i find it easier to find the vertex, but either way works